# Spectrum of Schrödinger operators on subcovering graphs

### v2

@@ 2.2. Spectra of Schrödinger operators on subcoverings with primitive sets of chiral vectors
see, e.g., [KK10]. The spectrum $\sigma(H_{\mathfrak{t}})$ consists of four spectral bands (since the fundamental graph of ${\bf G}_{2\mathfrak{a}_{1}}$ has four vertices, see Fig. 1e), two of which are flat bands.
ii) For any primitive $\mathfrak{t}\in\Gamma$, the spectrum of the Schrödinger operator $H_{\mathfrak{t}}$ on the nanotube ${\bf G}_{\mathfrak{t}}$ is symmetric with respect to the point 3 and has a gap of length $02(q+\varepsilon(\mathfrak{t}))>$ 0.
iii) The spectra of discrete (and differential) Schrödinger operators on nanotubes with any chiral vector $\mathfrak{t}$ (including non-primitive ones) were studied in [KP07]. Zig-zag ($t=(n,0)$) and armchair ($t=(n,n)$) nanotubes in magnetic and electric fields were considered in [KK10].
As the second example we consider the $d$-dimensional lattice ${\mathbb{L}}^{d}$, for $d=3$ see Fig. 2a. The lattice ${\mathbb{L}}^{d}$ is a ${\mathbb{Z}}^{d}$-periodic graph. The spectrum of the Laplacian $\Delta$ on ${\mathbb{L}}^{d}$ has the form $\sigma(\Delta)=[0,4d]$. We describe the Laplacian spectrum on the subcovering ${\mathbb{L}}^{d}_{\mathfrak{t}}$ of the lattice ${\mathbb{L}}^{d}$ with a primitive set $\mathfrak{t}$ of chiral vectors $\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\in{\mathbb{Z}}^{d}$, $d_{o}<d$ (for $d=3$, $d_{o}=1$ and $\mathfrak{t}_{1}=(1,1,-1)$ see Fig. 2b; for $d=3$, $d_{o}=2$ and $\mathfrak{t}_{1}=(1,1,-1)$, $\mathfrak{t}_{2}=(2,1,0)$ see Fig. 2c).
@@
${\mathcal{T}}$ is the chiral matrix of ${\mathbb{L}}^{d}_{\mathfrak{t}}$ defined by (1.9), and $\tau$ is the smallest eigenvalue of $({\mathcal{T}}{\mathcal{T}}^{\top})^{1/2}$. In particular, if $d_{o}=1$, then
$\lambda_{1}^{+}(\mathfrak{t})=4d-\dfrac{\pi^{2}}{\tau^{2}}+O\big{(}\frac{1}{\tau^{4}}\big{)},\qquad\textrm{as}\qquad\tau=|\mathfrak{t}_{1}|\to\infty.$ (2.11)
@@
ii) The triangular lattice shown in Fig. 2b is the subcovering ${\mathbb{L}}^{3}_{\mathfrak{t}_{1}}$ of the cubic lattice ${\mathbb{L}}^{3}$ with the chiral vector $\mathfrak{t}_{1}=(1,1,-1)$. It is known that the spectrum of the Laplacian $\Delta_{\mathfrak{t}_{1}}$ on the triangular lattice has the form $\sigma(\Delta_{\mathfrak{t}_{1}})=[0,9]$. On the other hand, since $\mathfrak{t}_{1}$ is primitive and the sum of its components is odd, then, due to Example 2.8, we have $\sigma(\Delta_{\mathfrak{t}_{1}})=[0,\lambda_{1}^{+}(\mathfrak{t}_{1})]$, and the asymptotics (2.11) (which is very rude in this case) yields
$\lambda_{1}^{+}(\mathfrak{t}_{1})\approx 4\cdot 3-\frac{\pi^{2}}{|\mathfrak{t}_{1}|^{2}}\approx 8{.}71.$
@@
After some simple calculations we obtain
@@
that agrees well with the spectrum $\sigma(\Delta_{\mathfrak{t}})=[0,11.34]$ calculated numerically. Note that the smallest eigenvalue of the matrix $({\mathcal{T}}{\mathcal{T}}^{\top})^{1/2}$ is $\tau\approx 3{.}42$.
Finally, we consider the diamond lattice ${\mathcal{G}}$ which is a $\Gamma$-periodic graph, where $\Gamma\subset{\mathbb{R}}^{3}$ is the lattice with the basis $\mathfrak{a}_{1},\mathfrak{a}_{2},\mathfrak{a}_{3}$, see Fig. 3a. The fundamental graph ${\mathcal{G}}_{*}$ of ${\mathcal{G}}$ consists of two vertices $v_{1}$ and $v_{2}$ and four multiple edges connecting these vertices, see Fig. 3d. Let $H=\Delta+Q$ be the Schrödinger operator on ${\mathcal{G}}$ with a $\Gamma$-periodic potential $Q$. Without loss of generality we may assume that the potential $Q$ is defined by
@@
Let ${\mathcal{G}}_{\mathfrak{t}}$ be a subcovering of the diamond lattice ${\mathcal{G}}$ with a primitive set $\mathfrak{t}\subset\Gamma$ of chiral vectors, and let $H_{\mathfrak{t}}=\Delta_{\mathfrak{t}}+Q$ be the Schrödinger operator on ${\mathcal{G}}_{\mathfrak{t}}$ with the periodic potential $Q$ defined by (2.12). Then $\sigma(H_{\mathfrak{t}})=\sigma(H)$ if and only if
$\bullet$ $\mathfrak{t}$ consists of a single primitive vector of the lattice $\Gamma$, or
$\bullet$ $\mathfrak{t}=\{\mathfrak{t}_{1},\mathfrak{t}_{2}\}$, where $\mathfrak{t}_{s}=t_{s1}\mathfrak{a}_{1}+t_{s2}\mathfrak{a}_{2}+t_{s3}\mathfrak{a}_{3}$, $s=1,2$, such that for some $i,j=1,2,3$, $i\neq j$, the sum $t_{si}+t_{sj}$ is even for each $s=1,2$.
@@
i) The subcovering ${\mathcal{G}}_{\mathfrak{t}}$ of the diamond lattice ${\mathcal{G}}$ with the primitive chiral vector $\mathfrak{t}=\mathfrak{a}_{1}+\mathfrak{a}_{2}-\mathfrak{a}_{3}$ shown in Fig. 3b is isospectral to ${\mathcal{G}}$.
@@
where $\varepsilon(\mathfrak{t})>0$ and $\varepsilon(\mathfrak{t})\to 0$ as $\tau\to\infty$. Recall that $\tau$ is the smallest eigenvalue of $({\mathcal{T}}{\mathcal{T}}^{\top})^{1/2}$, and ${\mathcal{T}}$ is the chiral matrix of ${\mathcal{G}}_{\mathfrak{t}}$ defined by (1.9). For example, if $\mathfrak{t}_{1}=\mathfrak{a}_{1}-\mathfrak{a}_{3}$ and $\mathfrak{t}_{2}=\mathfrak{a}_{2}-\mathfrak{a}_{3}$, then the spectrum of $H_{\mathfrak{t}}$ on the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ with the primitive set $\mathfrak{t}=\{\mathfrak{t}_{1},\mathfrak{t}_{2}\}$ shown in Fig. 3c is given by (2.14), where $\varepsilon(\mathfrak{t})=\sqrt{4+q^{2}}-q$.
@@ 3. Proof of the main results
Let ${\mathcal{G}}\subset{\mathbb{R}}^{d}$ be a $\Gamma$-periodic graph, where $\Gamma\subset{\mathbb{R}}^{d}$ is the lattice with the basis $\{\mathfrak{a}_{1},\ldots,\mathfrak{a}_{d}\}$. Recall that the band functions $\lambda_{j}(k)$ of the Schrödinger operator $H$ on ${\mathcal{G}}$ are $2\pi{\mathbb{Z}}^{d}$-periodic in $k\in{\mathbb{R}}^{d}$. In order to prove our main results we need a particular choice of the Brillouin zone for ${\mathcal{G}}$ instead of ${\mathcal{B}}=(-\pi,\pi]^{d}$. By a Brillouin zone we understand any choice of the fundamental domain of the lattice $2\pi{\mathbb{Z}}^{d}$ in the momentum space.
Let ${\mathcal{G}}_{\mathfrak{t}}$ be a subcovering of ${\mathcal{G}}$ with a primitive set $\mathfrak{t}=\{\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\}$, $d_{o}<d$, of the lattice $\Gamma$. This primitive set $\mathfrak{t}$ can be extended to a basis
@@
of $\Gamma$. The two bases of the same lattice $\Gamma$ are related by the $d\times d$ unimodular matrix $\widetilde{\mathcal{T}}$ (i.e., the integer matrix with the determinant ±$\pm 1$) given by
$\widetilde{\mathcal{T}}=(t_{ij})_{i,j=1}^{d},\quad\textrm{where}\quad\mathfrak{t}_{i}=\sum\limits_{j=1}^{d}t_{ij}\mathfrak{a}_{j},\qquad t_{ij}\in{\mathbb{Z}},\qquad\mathop{\mathrm{det}}\nolimits\widetilde{\mathcal{T}}=\pm 1.$ (3.2)
The rows of $\widetilde{\mathcal{T}}$ are the coordinates of $\mathfrak{t}_{s}\in\Gamma$, $s\in{\mathbb{N}}_{d}$, with respect to the basis $\{\mathfrak{a}_{1},\ldots,\mathfrak{a}_{d}\}$. Note that the first $d_{o}$ rows of $\widetilde{\mathcal{T}}$ form the chiral matrix ${\mathcal{T}}$ of the subcovering ${\mathcal{G}}_{\mathfrak{t}}$, see (1.8), (1.9). This change of the basis of the lattice $\Gamma$ determines the linear change in the quasimomentum $k$:
@@
Figure 4. Two choices of the Brillouin zone in terms of the quasimomenta $k=(k_{1},k_{2})\in(-\pi,\pi]^{2}$ (the square ${\mathcal{B}}$ bounded by dashed lines) and $\kappa=\widetilde{\mathcal{T}}k$, $\kappa=(\kappa_{1},\kappa_{2})\in(-\pi,\pi]^{2}$ (the shaded parallelogram $\widetilde{\mathcal{B}}$). The lines $L_{n}$ (marked in red color) are given by the equations $t_{1}k_{1}+t_{2}k_{2}=2\pi n$, $n\in{\mathbb{Z}}$, when $(t_{1},t_{2})=(2,3)$. The lines $L_{0},L_{\pm 1},L_{\pm 2}$ intersect the square Brillouin zone ${\mathcal{B}}$. The line $L_{0}$ (which coincides with the axis $\kappa_{2}$) is the single line among $L_{n}$ that enters the shaded Brillouin zone $\widetilde{\mathcal{B}}$.
@@
We consider the hexagonal lattice ${\bf G}$, which is a $\Gamma$-periodic graph, where $\Gamma\subset{\mathbb{R}}^{2}$ is the lattice with the basis $\{\mathfrak{a}_{1},\mathfrak{a}_{2}\}$, see Fig. 1a. Let ${\bf G}_{\mathfrak{t}_{1}}$ be the subcovering of ${\bf G}$ (the nanotube) with the primitive chiral vector $\mathfrak{t}_{1}=2\mathfrak{a}_{1}+3\mathfrak{a}_{2}$. The primitive vector $\mathfrak{t}_{1}$ can be completed to the basis $\{\mathfrak{t}_{1},\mathfrak{t}_{2}\}$ of $\Gamma$ by the vector $\mathfrak{t}_{2}=\mathfrak{a}_{1}+2\mathfrak{a}_{2}$. Thus, the matrix $\widetilde{\mathcal{T}}$ defined by (3.2) has the form
@@
$\bullet$ the square ${\mathcal{B}}$ (bounded by dashed lines) drawn in terms of the quasimomentum $k=(k_{1},k_{2})\in(-\pi,\pi]^{2}$, and
@@
The lines $L_{n}$ shown in the figure are given by the equations $t_{1}k_{1}+t_{2}k_{2}=2\pi n$, $n\in{\mathbb{Z}}$, where $(t_{1},t_{2})=(2,3)$ is the chiral index of the nanotube ${\bf G}_{\mathfrak{t}_{1}}$. The set ${\mathcal{B}}_{\mathfrak{t}_{1}}$ defined by (3.4) consists of the five segments of the lines $L_{0},L_{\pm 1},L_{\pm 2}$ inside the square Brillouin zone ${\mathcal{B}}$. The set ~$\widetilde{\mathcal{B}}_{\mathfrak{t}_{1}}$ defined by (3.6) consists of only the part of the line $L_{0}$ inside the Brillouin zone $\widetilde{\mathcal{B}}$. The line $L_{0}$ coincides with the axis $\kappa_{2}$.
The dispersion relation for the Schrödinger operator $H_{\bf G}=\Delta_{\bf G}+Q$ on ${\bf G}$ with a $\Gamma$-periodic potential $Q$ defined by (2.7) is plotted in Fig. 5 (left), when $q=1$. The Brillouin zone is parameterized by the quasimomentum $k$. The band functions of the Schrödinger operator $H_{{\bf G}_{\mathfrak{t}_{1}}}$ on the nanotube ${\bf G}_{\mathfrak{t}_{1}}$ are the restrictions of the dispersion relation for $H_{\bf G}$ to the segment of the line $L_{0}$ given by the equation $02k_{1}+3k_{2}=$ 0 (or, $\kappa_{1}=0$, in terms of $\kappa$) inside the Brillouin zone $\widetilde{\mathcal{B}}$, see Fig. 5 (right). The ranges of the band functions of $H_{{\bf G}}$ and $H_{{\bf G}_{\mathfrak{t}_{1}}}$ (i.e., the spectra of $H_{{\bf G}}$ and $H_{{\bf G}_{\mathfrak{t}_{1}}}$) have the form
$\sigma(H_{{\bf G}})=[3-\sqrt{10}\,,2]\cup[4,3+\sqrt{10}\,],\qquad\sigma(H_{{\bf G}_{\mathfrak{t}_{1}}})\approx[3-\sqrt{10}\,,1{.}92]\cup[4{.}08,3+\sqrt{10}\,].$
@@
Figure 5. Left: the graphs of the band functions (the dispersion relation) of the Schrödinger operator $H_{{\bf G}}$ (when $q=1$) on the hexagonal lattice ${\bf G}$. Right: the graphs of the band functions of the Schrödinger operator $H_{{\bf G}_{\mathfrak{t}_{1}}}$ on the nanotube ${\bf G}_{\mathfrak{t}_{1}}$ with the chiral vector $\mathfrak{t}_{1}=2\mathfrak{a}_{1}+3\mathfrak{a}_{2}$. The ranges of the corresponding band functions of $H_{{\bf G}}$ and $H_{{\bf G}_{\mathfrak{t}_{1}}}$ are close to each other but do not coincide.
@@
In fact, the set ~$\widetilde{\mathcal{B}}_{\mathfrak{t}}$ given by (3.6) is a natural choice of the Brillouin zone for the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ which is a $\Gamma_{o}$-periodic graph, where $\Gamma_{o}$ is the sublattice of $\Gamma$ of rank $d-d_{o}$ with the basis $\{\mathfrak{t}_{d_{o}+1},\ldots,\mathfrak{t}_{d}\}$, see (3.1).
@@
where $\tau$ is the smallest eigenvalue of $({\mathcal{T}}{\mathcal{T}}^{\top})^{1/2}$. Using the last inclusion in (2.2) and $2\pi{\mathbb{Z}}^{d}$-periodicity of the band function $\lambda_{j}(k)$, we have
@@
where ~$\widetilde{\mathcal{B}}_{\mathfrak{t}}$ is given by (3.6) and $L_{0}=\big{\{}k\in{\mathbb{R}}^{d}:{\mathcal{T}}k=0\big{\}}$. Thus, in order to get the band edge $\lambda^{-}_{j}(\mathfrak{t})$ we need to find the minimum of $\lambda_{j}(k)$ with the constraint ${\mathcal{T}}k=0$, where ${\mathcal{T}}$ is the chiral matrix of the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ defined by by (1.8), (1.9).
@@
Among all solutions of the system (3.13) we need to find the solution $\hat{y}$ with minimum Euclidean norm $|\hat{y}|$. The $d_{o}\times d$ matrix ${\mathcal{T}}{\bf H}^{-1/2}$ has full row rank. Thus, due to Proposition 3.4, the solution $\hat{y}$ of the system of the linear equations (3.13) with minimum $|\hat{y}|$ has the form
@@
Finally, let $d_{o}=1$, i.e., the set $\mathfrak{t}$ consists of a single primitive vector with a chiral index $t\in{\mathbb{Z}}^{d}$. Then ${\mathcal{T}}=t^{\top}$,
$\tau=(t^{\top}t)^{1/2}=|t|,\quad{\mathcal{T}}{\bf H}^{-1}{\mathcal{T}}^{\top}=t^{\top}{\bf H}^{-1}t=|{\bf H}^{-1/2}t|^{2},\quad{\bf x}_{o}=(\langle t,k_{o}\rangle\mod 2\pi)\in(-\pi,\pi],$

### v3

@@ 2.2. Spectra of Schrödinger operators on subcoverings with primitive sets of chiral vectors
see, e.g., [KK10]. The spectrum $\sigma(H_{\mathfrak{t}})$ consists of four spectral bands (since the fundamental graph of ${\bf G}_{\mathfrak{t}}$ has four vertices, see Fig. 1e), two of which are flat bands.
ii) For any primitive $\mathfrak{t}\in{\mathbb{Z}}^{2}$, the spectrum of the Schrödinger operator $H_{\mathfrak{t}}$ on the nanotube ${\bf G}_{\mathfrak{t}}$ is symmetric with respect to the point 3 and has a gap of length $02(q+\varepsilon(\mathfrak{t}))>$ 0.
iii) The spectra of discrete (and differential) Schrödinger operators on nanotubes with any chiral vector $\mathfrak{t}$ (including non-primitive ones) were studied in [KP07]. Zig-zag ($\mathfrak{t}=(n,0)$) and armchair ($\mathfrak{t}=(n,n)$) nanotubes in magnetic and electric fields were considered in [KK10].
As the second example we consider the $d$-dimensional lattice ${\mathbb{L}}^{d}$, for $d=3$ see Fig. 2a. The lattice ${\mathbb{L}}^{d}$ is a ${\mathbb{Z}}^{d}$-periodic graph with the periods $\mathfrak{a}_{1},\ldots,\mathfrak{a}_{d}$. The spectrum of the Laplacian $\Delta$ on ${\mathbb{L}}^{d}$ has the form $\sigma(\Delta)=[0,4d]$. We describe the Laplacian spectrum on the subcovering ${\mathbb{L}}^{d}_{\mathfrak{t}}$ of the lattice ${\mathbb{L}}^{d}$ with a primitive set $\mathfrak{t}$ of chiral vectors $\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\in{\mathbb{Z}}^{d}$, $d_{o}<d$ (for $d=3$, $d_{o}=1$ and $\mathfrak{t}_{1}=(1,1,-1)$ see Fig. 2b; for $d=3$, $d_{o}=2$ and $\mathfrak{t}_{1}=(1,1,-1)$, $\mathfrak{t}_{2}=(2,1,0)$ see Fig. 2c).
@@
${\mathcal{T}}$ is the chiral matrix of ${\mathbb{L}}^{d}_{\mathfrak{t}}$ defined by (1.8), and $\tau$ is the smallest singular value of ${\mathcal{T}}$.
In particular, if ${\mathbb{L}}^{d}_{\mathfrak{t}}$ is a subcovering of ${\mathbb{L}}^{d}$ with a single primitive vector $\mathfrak{t}=(t_{1},\ldots,t_{d})\in{\mathbb{Z}}^{d}$ and $t_{1}+\ldots+t_{d}$ is odd, then
$\lambda_{1}^{+}(\mathfrak{t})=4d-\dfrac{\pi^{2}}{\tau^{2}}+O\big{(}\frac{1}{\tau^{4}}\big{)},\qquad\textrm{as}\qquad\tau=|\mathfrak{t}|\to\infty.$ (2.12)
@@
ii) The triangular lattice shown in Fig. 2b is the subcovering ${\mathbb{L}}^{3}_{\mathfrak{t}}$ of the cubic lattice ${\mathbb{L}}^{3}$ with the chiral vector $\mathfrak{t}=(1,1,-1)$. It is known that the spectrum of the Laplacian $\Delta_{\mathfrak{t}}$ on the triangular lattice has the form $\sigma(\Delta_{\mathfrak{t}})=[0,9]$. On the other hand, since $\mathfrak{t}$ is primitive and the sum of its components is odd, then, due to Example 2.8, we have $\sigma(\Delta_{\mathfrak{t}})=[0,\lambda_{1}^{+}(\mathfrak{t})]$, and the asymptotics (2.12) (which is very rude in this case) yields
$\lambda_{1}^{+}(\mathfrak{t})\approx 4\cdot 3-\frac{\pi^{2}}{|\mathfrak{t}|^{2}}\approx 8{.}71.$
@@
After some simple calculations, we obtain
@@
that agrees well with the spectrum $\sigma(\Delta_{\mathfrak{t}})=[0,11.34]$ calculated numerically. Note that the smallest singular value of the matrix ${\mathcal{T}}$ is $\tau\approx 3{.}42$.
Finally, we consider the diamond lattice ${\mathcal{G}}$ which is a ${\mathbb{Z}}^{3}$-periodic graph with the periods $\mathfrak{a}_{1},\mathfrak{a}_{2},\mathfrak{a}_{3}$, see Fig. 3a. The fundamental graph ${\mathcal{G}}_{*}$ of ${\mathcal{G}}$ consists of two vertices $v_{1}$ and $v_{2}$ and four multiple edges connecting these vertices, see Fig. 3d. Let $H=\Delta+Q$ be the Schrödinger operator on ${\mathcal{G}}$ with a ${\mathbb{Z}}^{3}$-periodic potential $Q$. Without loss of generality we may assume that the potential $Q$ is defined by
@@
Let ${\mathcal{G}}_{\mathfrak{t}}$ be a subcovering of the diamond lattice ${\mathcal{G}}$ with a primitive set $\mathfrak{t}\subset{\mathbb{Z}}^{3}$ of chiral vectors, and let $H_{\mathfrak{t}}=\Delta_{\mathfrak{t}}+Q$ be the Schrödinger operator on ${\mathcal{G}}_{\mathfrak{t}}$ with the periodic potential $Q$ defined by (2.13). Then $\sigma(H_{\mathfrak{t}})=\sigma(H)$ if and only if
$\bullet$ $\mathfrak{t}$ consists of a single primitive vector, or
$\bullet$ $\mathfrak{t}=\{\mathfrak{t}_{1},\mathfrak{t}_{2}\}$, where $\mathfrak{t}_{s}=(t_{s1},t_{s2},t_{s3})\in{\mathbb{Z}}^{3}$, $s=1,2$, such that for some $i,j=1,2,3$, $i\neq j$, the sum $t_{si}+t_{sj}$ is even for each $s=1,2$.
@@
i) The subcovering ${\mathcal{G}}_{\mathfrak{t}}$ of the diamond lattice ${\mathcal{G}}$ with the primitive chiral vector $\mathfrak{t}=(1,1,-1)$ shown in Fig. 3b is isospectral to ${\mathcal{G}}$.
@@
where $\varepsilon(\mathfrak{t})>0$ and $\varepsilon(\mathfrak{t})\to 0$ as $\tau\to\infty$. Recall that $\tau$ is the smallest singular value of the chiral matrix ${\mathcal{T}}$ of ${\mathcal{G}}_{\mathfrak{t}}$ defined by (1.8). For example, if $\mathfrak{t}_{1}=(1,0,-1)$ and $\mathfrak{t}_{2}=(0,1,-1)$, then the spectrum of $H_{\mathfrak{t}}$ on the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ with the primitive set $\mathfrak{t}=\{\mathfrak{t}_{1},\mathfrak{t}_{2}\}$ shown in Fig. 3c is given by (2.15), where $\varepsilon(\mathfrak{t})=\sqrt{4+q^{2}}-q$.
@@ 3. Proof of the main results
Let ${\mathcal{G}}$ be a ${\mathbb{Z}}^{d}$-periodic graph with the periods $\mathfrak{a}_{1},\ldots,\mathfrak{a}_{d}$. Recall that the band functions $\lambda_{j}(k)$ of the Schrödinger operator $H$ on ${\mathcal{G}}$ are $2\pi{\mathbb{Z}}^{d}$-periodic in $k\in{\mathbb{R}}^{d}$. In order to prove our main results we need a particular choice of the Brillouin zone for ${\mathcal{G}}$ instead of ${\mathcal{B}}=(-\pi,\pi]^{d}$. By a Brillouin zone we understand any choice of the fundamental domain of the lattice $2\pi{\mathbb{Z}}^{d}$ in the momentum space.
Let ${\mathcal{G}}_{\mathfrak{t}}$ be a subcovering of ${\mathcal{G}}$ with a primitive set $\mathfrak{t}=\{\mathfrak{t}_{1},\ldots,\mathfrak{t}_{d_{o}}\}\subset{\mathbb{Z}}^{d}$, $d_{o}<d$. This primitive set $\mathfrak{t}$ can be extended to a basis
@@
of ${\mathbb{Z}}^{d}$. We denote by $\widetilde{\mathcal{T}}$ the $d\times d$ unimodular matrix (i.e., the integer matrix with the determinant ±$\pm 1$) whose rows are the vectors $\mathfrak{t}_{1},\ldots\mathfrak{t}_{d}$:
$\widetilde{\mathcal{T}}=(t_{ij})_{i,j=1}^{d},\quad\textrm{where}\quad\mathfrak{t}_{i}=(t_{i1},\ldots,t_{id})\in{\mathbb{Z}}^{d},\quad i\in{\mathbb{N}}_{d},\quad\mathop{\mathrm{det}}\nolimits\widetilde{\mathcal{T}}=\pm 1.$ (3.2)
Note that the first $d_{o}$ rows of $\widetilde{\mathcal{T}}$ form the chiral matrix ${\mathcal{T}}$ of the subcovering ${\mathcal{G}}_{\mathfrak{t}}$, see (1.8). The change of the basis of ${\mathbb{Z}}^{d}$ (from $\{\mathfrak{a}_{1},\ldots,\mathfrak{a}_{d}\}$ to the basis (3.1)) determines the linear change in the quasimomentum $k$:
@@
Figure 4. Two choices of the Brillouin zone in terms of the quasimomenta $k=(k_{1},k_{2})\in(-\pi,\pi]^{2}$ (the square ${\mathcal{B}}$ bounded by the dashed lines) and $\kappa=\widetilde{\mathcal{T}}k$, $\kappa=(\kappa_{1},\kappa_{2})\in(-\pi,\pi]^{2}$ (the shaded parallelogram $\widetilde{\mathcal{B}}$). The lines $L_{n}$ (marked in red color) are given by the equations $t_{1}k_{1}+t_{2}k_{2}=2\pi n$, $n\in{\mathbb{Z}}$, when $(t_{1},t_{2})=(2,3)$. The lines $L_{0},L_{\pm 1},L_{\pm 2}$ intersect the square Brillouin zone ${\mathcal{B}}$. The line $L_{0}$ (which coincides with the axis $\kappa_{2}$) is the single line among $L_{n}$ that enters the shaded Brillouin zone $\widetilde{\mathcal{B}}$.
@@
We consider the hexagonal lattice ${\bf G}$, which is a ${\mathbb{Z}}^{2}$-periodic graph with the periods $\mathfrak{a}_{1},\mathfrak{a}_{2}$, see Fig. 1a. Let ${\bf G}_{\mathfrak{t}_{1}}$ be the subcovering of ${\bf G}$ (the nanotube) with the primitive chiral vector $\mathfrak{t}_{1}=(2,3)$. The primitive vector $\mathfrak{t}_{1}=(2,3)$ can be completed to the basis $\{\mathfrak{t}_{1},\mathfrak{t}_{2}\}$ of ${\mathbb{Z}}^{2}$ by the vector $\mathfrak{t}_{2}=(1,2)$. Thus, the matrix $\widetilde{\mathcal{T}}$ defined by (3.2) has the form
@@
$\bullet$ the square ${\mathcal{B}}$ (bounded by the dashed lines) drawn in terms of the quasimomentum $k=(k_{1},k_{2})\in(-\pi,\pi]^{2}$, and
@@
The lines $L_{n}$ shown in the figure are given by the equations $t_{1}k_{1}+t_{2}k_{2}=2\pi n$, $n\in{\mathbb{Z}}$, where $\mathfrak{t}_{1}=(t_{1},t_{2})=(2,3)$ is the chiral vector of the nanotube ${\bf G}_{\mathfrak{t}_{1}}$. The set ${\mathcal{B}}_{\mathfrak{t}_{1}}$ defined by (3.4) consists of the five segments of the lines $L_{0},L_{\pm 1},L_{\pm 2}$ inside the square Brillouin zone ${\mathcal{B}}$. The set ~$\widetilde{\mathcal{B}}_{\mathfrak{t}_{1}}$ defined by (3.6) consists of only the part of the line $L_{0}$ inside the Brillouin zone $\widetilde{\mathcal{B}}$. The line $L_{0}$ coincides with the axis $\kappa_{2}$.
The dispersion relation for the Schrödinger operator $H=\Delta+Q$ on ${\bf G}$ with the unit edge weight and a ${\mathbb{Z}}^{2}$-periodic potential $Q$ defined by (2.8) is plotted in Fig. 5 (left and center), when $q=1$. The Brillouin zone is parameterized by the quasimomentum $k$ in the left figure and by the quasimomentum $\kappa=\widetilde{\mathcal{T}}k$ in the center figure. The band functions of the Schrödinger operator $H_{\mathfrak{t}_{1}}$ on the nanotube ${\bf G}_{\mathfrak{t}_{1}}$ are the restrictions of the dispersion relation for $H$ to the segment of the line $L_{0}$ given by the equation $02k_{1}+3k_{2}=$ 0 (or, $\kappa_{1}=0$, in terms of $\kappa$) inside the Brillouin zone $\widetilde{\mathcal{B}}$, see Fig. 5 (right). The ranges of the band functions of $H$ and $H_{\mathfrak{t}_{1}}$ (i.e., the spectra of $H$ and $H_{\mathfrak{t}_{1}}$) have the form
$\sigma(H)=[3-\sqrt{10}\,,2]\cup[4,3+\sqrt{10}\,],\qquad\sigma(H_{\mathfrak{t}_{1}})\approx[3-\sqrt{10}\,,1{.}92]\cup[4{.}08,3+\sqrt{10}\,].$
@@
Figure 5. Left and center: the graphs of the band functions (the dispersion relation) of the Schrödinger operator $H$ (when $q=1$) on the hexagonal lattice ${\bf G}$. The Brillouin zone is parameterized by the quasimomenta $k=(k_{1},k_{2})$ (left) and $\kappa=(\kappa_{1},\kappa_{2})$, $\kappa=\widetilde{\mathcal{T}}k$ (center). Right: the graphs of the band functions of the Schrödinger operator $H_{\mathfrak{t}_{1}}$ on the nanotube ${\bf G}_{\mathfrak{t}_{1}}$ with the chiral vector $\mathfrak{t}_{1}=(2,3)$. The band functions of $H_{\mathfrak{t}_{1}}$ are the restrictions of the dispersion relation for $H$ (center) to the axis $\kappa_{2}$. The ranges of the corresponding band functions of $H$ and $H_{\mathfrak{t}_{1}}$ are close to each other but do not coincide.
@@
In fact, the set ~$\widetilde{\mathcal{B}}_{\mathfrak{t}}$ given by (3.6) is a natural choice of the Brillouin zone for the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ which is a ${\mathbb{Z}}^{d-d_{o}}$-periodic graph with the periods $\mathfrak{t}_{d_{o}+1},\ldots,\mathfrak{t}_{d}$, see (3.1).
@@
where $\tau$ is the smallest singular value of ${\mathcal{T}}$. Using the last inclusion in (2.2) and $2\pi{\mathbb{Z}}^{d}$-periodicity of the band function $\lambda_{j}(k)$, we have
@@
where ~$\widetilde{\mathcal{B}}_{\mathfrak{t}}$ is given by (3.6) and $L_{0}=\big{\{}k\in{\mathbb{R}}^{d}:{\mathcal{T}}k=0\big{\}}$. Thus, in order to get the band edge $\lambda^{-}_{j}(\mathfrak{t})$ we need to find the minimum of $\lambda_{j}(k)$ with the constraint ${\mathcal{T}}k=0$, where ${\mathcal{T}}$ is the chiral matrix of the subcovering ${\mathcal{G}}_{\mathfrak{t}}$ defined by (1.8).
@@
Among all solutions of the system (3.13) we need to find the solution $\hat{y}$ with minimum Euclidean norm $|\hat{y}|$. The $d_{o}\times d$ matrix ${\mathcal{T}}{\bf H}^{-1/2}$ has full row rank. Thus, due to Proposition 3.4, the solution $\hat{y}$ of the system of linear equations (3.13) with minimum $|\hat{y}|$ has the form
@@
Finally, let ${\mathcal{G}}_{\mathfrak{t}}$ be a subcovering of ${\mathcal{G}}$ with a single primitive vector $\mathfrak{t}\in{\mathbb{Z}}^{d}$. Then ${\mathcal{T}}=\mathfrak{t}^{\top}$,
$\tau=(\mathfrak{t}^{\top}\mathfrak{t})^{1/2}=|\mathfrak{t}|,\quad{\mathcal{T}}{\bf H}^{-1}{\mathcal{T}}^{\top}=\mathfrak{t}^{\top}{\bf H}^{-1}\mathfrak{t}=|{\bf H}^{-1/2}\mathfrak{t}|^{2},\quad{\bf x}_{o}=(\langle\mathfrak{t},k_{o}\rangle\mod 2\pi)\in(-\pi,\pi],$